# Off-policy Evaluation with Deeply-abstracted States

### v2

@@
Abstract. Off-policy evaluation (OPE) is crucial for assessing a target policy’s impact offline before its deployment. However, achieving accurate OPE in large state spaces remains challenging. This paper studies state abstractions – originally designed for policy learning – in the context of OPE. Our contributions are three-fold: (i) We define a set of irrelevance conditions central to learning state abstractions for OPE, and derive a backward-model-irrelevance condition for achieving irrelevance in (marginalized) importance sampling ratios by constructing a time-reversed Markov decision process (MDP). (ii) We propose a novel iterative procedure that sequentially projects the original state space into a smaller space, resulting in a deeply-abstracted state, which substantially simplifies the sample complexity of OPE arising from high cardinality. (iii) We prove the Fisher consistencies of various OPE estimators when applied to our proposed abstract state spaces.
@@ 1 Introduction
Motivation. Off-policy evaluation (OPE) serves as a crucial tool for assessing the impact of a newly developed policy using a pre-collected historical data before its deployment in high-stake applications, such as healthcare (Murphy et al.,, 2001), recommendation systems (Chapelle and Li,, 2011), education (Mandel et al.,, 2014), dialog systems (Jiang et al.,, 2021) and robotics (Levine et al.,, 2020). A fundamental challenge in OPE is its “off-policy” nature, wherein the target policy to be evaluated differs from the behavior policy that generates the offline data. This distributional shift is particularly pronounced in environments with large state spaces of high cardinality. Theoretically, the error bounds for estimating the target policy’s Q-function and value decrease rapidly as the state space dimension increases (Hao et al., 2021a, ; Chen and Qi,, 2022). Empirically, large state space significantly challenges the performance of state-of-the-art OPE algorithms (Fu et al.,, 2020; Voloshin et al.,, 2021).
Although different policies induce different trajectories in the large ground state space, they can produce similar paths when restricted to relevant, lower-dimensional state spaces (Pavse and Hanna,, 2023). Consequently, applying OPE to these abstract spaces can significantly mitigate the distributional shift between target and behavior policies, enhancing the accuracy in predicting the target policy’s value. This makes state abstraction, designed to reduce state space cardinality, particularly appealing for OPE. However, despite the extensive literature on studying state abstractions for policy learning (see Section 2 for details), it has been hardly explored in the context of OPE.
@@
    Introduction of a set of irrelevance conditions for OPE, and derivation of a backward-model-irrelevance condition for state abstractions to achieve irrelevance in marginalized importance sampling ratios by constructing a time-reversed Markov decision process (MDP, Puterman,, 2014) that swaps the future and past.
@@
    Development of a novel iterative procedure to sequentially compress the state space. Specifically, within each iteration, our algorithm consistently produces a state space that is either smaller in size or remains the same. Through its iterative nature, the proposed approach produces a deeply-abstracted state space, which substantially reduces the sample complexity of OPE.
@@
    Validations of various OPE methods when applied to the proposed abstract state spaces.
Organization. The rest of the paper is structured as follows. Section 2 is dedicated to the literature review of related works. MDP-related notions and OPE methodologies relevant to our proposal are recalled in Section 3. Our proposed state abstractions for OPE are presented in Section 4. Section 5 conducts numerical experiments to demonstrate the efficiency of our approach.
@@ 2 Related work
Off-policy evaluation. OPE aims to estimate the expected return of a given target policy, utilizing historical data generated by a possibly different behavior policy (Dudík et al.,, 2014; Uehara et al.,, 2022). The majority of methods in the literature can be classified into the following three categories:
@@
    Value-based methods that estimate the target policy’s return by learning either a value function (Sutton et al.,, 2008; Luckett et al.,, 2019; Li et al.,, 2024) or a Q-function (Le et al.,, 2019; Feng et al.,, 2020; Hao et al., 2021b, ; Liao et al.,, 2021; Chen and Qi,, 2022; Shi et al.,, 2022) from the data.
@@
    Importance sampling (IS) methods that adjust the observed rewards using the IS ratio, i.e., the ratio of the target policy over the behavior policy, to address their distributional shift. There are two major types: sequential IS (SIS, Precup,, 2000; Thomas et al.,, 2015; Hanna et al.,, 2019; Hu and Wager,, 2023) which employs a cumulative IS ratio, and marginalized IS (MIS, Liu et al.,, 2018; Nachum et al.,, 2019; Xie et al.,, 2019; Dai et al.,, 2020; Yin and Wang,, 2020; Wang et al.,, 2023) which uses the MIS ratio to mitigate the high variance of the SIS estimator.
@@
    Doubly robust methods or their variants that employ both the IS ratio and the value/reward function to enhance the robustness of OPE (Zhang et al.,, 2013; Jiang and Li,, 2016; Thomas and Brunskill,, 2016; Farajtabar et al.,, 2018; Kallus and Uehara,, 2020; Tang et al.,, 2020; Uehara et al.,, 2020; Shi et al.,, 2021; Kallus and Uehara,, 2022; Liao et al.,, 2022; Xie et al.,, 2023).
@@
State abstraction. State abstraction aims to obtain a parsimonious state representation to simplify the sample complexity of reinforcement learning (RL), while ensuring that the optimal policy restricted to the abstract state space attains comparable values as in the original, ground state space. There is an extensive literature on the theoretical and methodological development of state abstraction, particularly bisimulation — a type of abstractions that preserve the Markov property in the abstracted state (Singh et al.,, 1994; Dean and Givan,, 1997; Givan et al.,, 2003; Ferns et al.,, 2004; Ravindran,, 2004; Jong and Stone,, 2005; Li et al.,, 2006; Ferns et al.,, 2011; Abel et al.,, 2016; Wang et al.,, 2017; Castro,, 2020; Allen et al.,, 2021; Abel,, 2022). In particular, Li et al., (2006) analyzed five irrelevance conditions for optimal policy learning. Unlike the aforementioned works that focus on policy learning, we introduce irrelevance conditions for OPE, and propose abstractions that satisfy these irrelevant properties. Meanwhile, the proposed abstraction for achieving irrelevance for the MIS ratio resembles the Markov state abstraction developed by Allen et al., (2021) in the context of policy learning, while relaxing their requirement for the behavior policy to be Markovian.
More recently, Pavse and Hanna, (2023) made a pioneering attempt to study state abstraction for OPE, proving its benefits in enhancing OPE accuracy. However, they primarily focused on MIS estimators. In contrast, our theoretical analysis applies to a broader range of OPE estimators, covering all three aforementioned categories. Moreover, their abstraction did not achieve MIS-ratio irrelevance. Nor did they implement the iterative procedure.
Lastly, state abstraction is also related to variable selection (Kolter and Ng,, 2009; Geist and Scherrer,, 2011; Geist et al.,, 2012; Nguyen et al.,, 2013; Fan et al.,, 2016; Guo and Brunskill,, 2017; Shi et al.,, 2018; Zhang and Zhang,, 2018; Qi et al.,, 2020; Hao et al., 2021a, ; Ma et al.,, 2023) as well as representation learning for both policy learning (see e.g., Gelada et al.,, 2019; Zhang et al.,, 2020; Uehara et al., 2021b, ) and OPE (see e.g., Wang et al.,, 2021; Chang et al.,, 2022; Ni et al.,, 2023; Pavse and Hanna,, 2024).
@@ 3 Preliminaries
In this section, we first introduce some key concepts relevant to OPE in RL, such as MDP, target and behavior policies, value functions, IS ratios (Section 3.1). We next review state abstractions for optimal policy learning (Section 3.2), alongside with four prominent OPE methodologies (Section 3.3).
@@ 3.1 Data generating process, policy, value and IS ratio
Data. Assume the offline dataset $\mathcal{D}$ comprises multiple trajectories, each containing a sequence of state-action-reward triplets $(S_{t},A_{t},R_{t})_{t\geq 1}$ following a finite MDP, denoted by $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\rho_{0},\gamma\rangle$. Here, $\mathcal{S}$ and $\mathcal{A}$ are the discrete state and action spaces, both with finite cardinalities, $\mathcal{T}$ and $\mathcal{R}$ are the state transition and reward functions, $\rho_{0}$ denotes the initial state distribution, and $\gamma\in(0,1)$ is the discount factor. The data is generated as follows:
@@
    The environment delivers an immediate reward $R_{t}$ with an expected value of $\mathcal{R}(A_{t},S_{t})$, and transits into the next state $S_{t+1}\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{T}(\bullet\mid A_{t},S_{t})$ according to the transition function $\mathcal{T}$.
@@
These functions are pivotal in developing value-based estimators, as described in Method 1 of Section 3.3. Moreover, we use $\pi^{*}$ to denote the optimal policy that maximizes $J(\pi)$, i.e., $\pi^{*}\in\arg\max_{\pi}J(\pi)$, and write the optimal Q- and value functions $Q^{\pi^{*}}$, $V^{\pi^{*}}$ as $Q^{*}$, $V^{*}$ for brevity.

### v3

@@
Abstract
Off-policy evaluation (OPE) is crucial for assessing a target policy’s impact offline before its deployment. However, achieving accurate OPE in large state spaces remains challenging. This paper studies state abstractions – originally designed for policy learning – in the context of OPE. Our contributions are three-fold: (i) We derive a bias-variance decomposition for the mean squared error (MSE) of various post-abstraction OPE estimators, e.g., estimators when applied to the abstract state space. Based on this decomposition, we demonstrate that the iterative application of Markov state abstractions (MSAs) can sequentially reduce the MSE of the resulting OPE estimator, ultimately leading to a substantial error reduction. In general, the finer the abstract state space, the lower the MSE of these estimators. (ii) Building on these insights, we propose a novel iterative procedure that sequentially projects the original state space into a smaller space while preserving the Markov assumption. This results in a deeply-abstracted state space, which is substantially smaller than the original and considerably reduces the sample complexity of OPE. (iii) We conduct numerical experiments to demonstrate the advantages of the proposed state abstraction for OPE in finite samples and to validate our theoretical findings.
@@ 1 Introduction
Motivation. Off-policy evaluation (OPE) serves as a crucial tool for assessing the impact of a newly developed policy using a pre-collected dataset before its deployment in high-stake applications, such as healthcare (Murphy et al., 2001), recommendation systems (Chapelle and Li, 2011), education (Mandel et al., 2014), dialog systems (Jiang et al., 2021) and robotics (Levine et al., 2020). A fundamental challenge in OPE is its “off-policy” nature, wherein the target policy to be evaluated differs from the behavior policy that generates the offline data. This distributional shift is particularly pronounced in environments with large state spaces of high cardinality. Theoretically, the error bounds for estimating the target policy’s Q-function and value decrease rapidly as the state space dimension increases (Hao et al., 2021a; Chen and Qi, 2022). Empirically, large state space significantly challenges the performance of state-of-the-art OPE algorithms (Fu et al., 2020; Voloshin et al., 2021).
Although different policies induce different trajectories in the large ground state space, they can produce similar paths when restricted to relevant, lower-dimensional state spaces (Pavse and Hanna, 2023). Consequently, applying OPE to these abstract spaces can substantially mitigate the distributional shift between target and behavior policies, enhancing the accuracy in predicting the target policy’s value. This makes state abstraction, designed to reduce state space cardinality, particularly appealing for OPE. However, despite the extensive literature on studying state abstractions for policy learning (see Section 2 for details), it has been hardly explored in the context of OPE.
@@
    Derivation of a bias-variance decomposition for the mean squared error (MSE) of various OPE estimators. Based on this decomposition, we demonstrate that the iterative application of Markov state abstractions (MSAs) can sequentially reduce the MSE of the resulting OPE estimator and ultimately lead to a substantial reduction in MSE. In general, the finer the abstract state space, the lower the MSE of these estimators.
@@
    Development of a novel iterative state abstraction procedure sequentially compresses the state space. Specifically, within each iteration, our algorithm projects the original state space into a smaller space while preserving the Markov assumption. This results in a deeply-abstracted state, which aims to maximally compress the state space and reduce the sample complexity of OPE in large state spaces.
@@
    Empirical validations of OPE methods when applied to the proposed abstract state spaces. Our results highlight (i) the effectiveness of state abstraction in improving OPE, (ii) the benefits of multi-iteration abstractions over single-iteration abstractions and (iii) the superiority of the proposed state abstraction over existing approaches.
Organization. The rest of the paper is structured as follows. Section 2 is dedicated to the literature review of related works. MDP-related notions and OPE methodologies relevant to our proposal are recalled in Section 3. Our unified theory for the MSE decomposition and proposed state abstractions for OPE are presented in Section 4. Section 5 conducts numerical experiments to demonstrate the efficiency of our approach and to validate our theories. Finally, Section 6 concludes the paper.
@@ 2 Related work
Off-policy evaluation. OPE aims to estimate the expected return of a given target policy, utilizing historical data generated by a possibly different behavior policy (Dudík et al., 2014; Uehara et al., 2022). The majority of methods in the literature can be classified into the following three categories:
@@
    Value-based methods that estimate the target policy’s return by learning either a value function (Sutton et al., 2008; Luckett et al., 2019; Li et al., 2024) or a Q-function (Le et al., 2019; Feng et al., 2020; Hao et al., 2021b; Liao et al., 2021; Chen and Qi, 2022; Shi et al., 2022) from the data.
@@
    Importance sampling (IS) methods that adjust the observed rewards using the IS ratio, i.e., the ratio of the target policy over the behavior policy, to address their distributional shift. There are two major types: sequential IS (SIS, Precup, 2000; Thomas et al., 2015; Hanna et al., 2019; Hu and Wager, 2023) which employs the product of IS ratios, and marginalized IS (MIS, Liu et al., 2018; Nachum et al., 2019; Xie et al., 2019; Dai et al., 2020; Yin and Wang, 2020; Wang et al., 2023) which uses the MIS ratio to mitigate the high variance of the SIS estimator.
@@
    Doubly robust methods or their variants that employ both the IS ratio and the value/reward function to enhance the robustness of OPE (Zhang et al., 2013; Jiang and Li, 2016; Thomas and Brunskill, 2016; Farajtabar et al., 2018; Kallus and Uehara, 2020; Tang et al., 2020; Uehara et al., 2020; Shi et al., 2021; Kallus and Uehara, 2022; Liao et al., 2022; Xie et al., 2023).
@@
State abstraction. State abstraction aims to obtain a parsimonious state representation to simplify the sample complexity of reinforcement learning (RL), while ensuring that the optimal policy restricted to the abstract state space attains comparable values as in the original, ground state space. There is an extensive literature on the theoretical and methodological development of state abstraction, particularly bisimulation — a type of abstractions that preserve the Markov property (Singh et al., 1994; Dean and Givan, 1997; Givan et al., 2003; Ferns et al., 2004; Ravindran, 2004; Jong and Stone, 2005; Li et al., 2006; Ferns et al., 2011; Abel et al., 2016; Wang et al., 2017; Castro, 2020; Allen et al., 2021; Abel, 2022). In particular, Li et al. (2006) analyzed five irrelevance conditions for optimal policy learning. Unlike the aforementioned works that focus on policy learning, we focus on OPE. Meanwhile, the proposed abstraction is closely related the Markov state abstraction developed by Allen et al. (2021) in the context of policy learning, while relaxing their requirement for the behavior policy to be Markovian.
More recently, Pavse and Hanna (2023) made a pioneering attempt to study state abstraction for OPE, proving its benefits in improving the accuracy of MIS estimators. Pavse and Hanna (2024) introduced a new diffuse metric to measure the similarity between state-action pairs and employed it to learn state-action representations for subsequent use in fitted Q-evaluation (FQE, Le et al., 2019) – a popular value-based OPE method. Different from these methods that are tailored to either MIS or FQE, our analysis applies to a broader range of OPE estimators, covering all three aforementioned categories. Chaudhari et al. (2024) leveraged state abstraction to distill complex, continuous systems into compact, discrete models for OPE. They theoretically showed that certain abstractions yield consistent OPE estimators applied to the abstract space. However, they did not theoretically demonstrate that state abstraction can reduce the MSE of the resulting OPE estimator, nor did they provide a comprehensive discussion on the approaches for computing these abstractions.
Lastly, state abstraction is also related to variable selection (Kolter and Ng, 2009; Geist and Scherrer, 2011; Geist et al., 2012; Nguyen et al., 2013; Fan et al., 2016; Guo and Brunskill, 2017; Shi et al., 2018; Zhang and Zhang, 2018; Qi et al., 2020; Hao et al., 2021a; Ma et al., 2023) as well as representation learning for both policy learning (see e.g., Gelada et al., 2019; Zhang et al., 2020; Uehara et al., 2021) and OPE (see e.g., Wang et al., 2021; Chang et al., 2022; Ni et al., 2023).
We remark that none of the aforementioned works considered the iterative use of state abstraction for accurate OPE, which is the cornerstone of our methodology.
@@ 3 Preliminaries
In this section, we first introduce some key concepts relevant to OPE in RL, such as MDP, target and behavior policies, value functions, IS ratios (Section 3.1). Subsequently, we review four prominent OPE methodologies in Section 3.2.
@@ 3.1 Data generating process, policy, value and IS ratio
Data. Assume the offline dataset $\mathcal{D}$ comprises $N$ trajectories, each containing a sequence of state-action-reward-next-state tuples $(S_{t},A_{t},R_{t},S_{t+1})_{1\leq t\leq T}$ following a finite MDP, denoted by $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\rho_{0},\gamma\rangle$. Here, $\mathcal{S}$ and $\mathcal{A}$ are the discrete state and action spaces, both with finite cardinalities, $\mathcal{T}$ and $\mathcal{R}$ are the state transition and reward functions, $\rho_{0}$ denotes the initial state distribution, and $\gamma\in(0,1)$ is the discount factor. The data is generated as follows:
@@
    The environment delivers an immediate reward $R$ with an expected value of $\mathcal{R}(A_{t},S_{t})$, and transits into the next state $S_{t+1}\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{T}(\bullet\mid A_{t},S_{t})$ according to the transition function $\mathcal{T}$.
@@
These functions are pivotal in developing value-based estimators, as described in Method 1 of Section 3.2.